Reid is 5 years older than Peter. In 9 years, the sum of their ages will be 57 years. How old is each now?
Answer by TimothyLamb(4379)

p + 5 = r
(p+9) + (r+9) = 57
---
p + r + 18 = 57
p + (p + 5) + 18 = 57
2p + 23 = 57
2p= 34
Answer:
p = 17
r = 22
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php